You have looked at the current financial statements for Reigle Homes, Co. The company has an EBIT of $3,190,000 this year. Depreciation, the increase in net working capital, and capital spending were $242,000, $107,000, and $500,000, respectively. You expect that over the next five years, EBIT will grow at 16 percent per year, depreciation and capital spending will grow at 21 per year, and NWC will grow at 11 per year. The company currently has $18,500,000 in debt and 385,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 3.5 percent indefinitely. The company’s WACC is 9.1 percent and the tax rate is 40 percent.

What is the price per share of the company's stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Share price            $
Answer #1

SOLUTION:-B E 1 Year 3 4,979,258.2400 1.991.703.2960 2.987,554.9440 428.717.7620 14,501.8170 885.780.5000 2.515,990.3890 А C D Particul

B Particulars Year 0 Yearl 2 Earnings before tax and interest 3190000 =B2*(1-16%) 3 Less: Tax a 40% =B2*40% =C2*40% 4 Net inc

Therefore, the price per share is $80.69